# Application of Quantum Annealing to Computation of Molecular Properties

### v1

@@ Abstract
In summary, we have demonstrated the efficacy of QAE algorithm in computing the PDMs of molecules. Unlike the traditional applications of QAE, which focus on finding the ground-state energy, our approach integrates perturbative techniques and the FFM to enable the calculation of molecular properties. This novel framework shows promise for a broad range of many-body physics applications. Key contributions of our work include: (i) The integration of the finite-field approach into the QAE workflow, (ii) An improved optimization strategy for the Lagrange multipliers within the energy functional, (iii) Perturbative selection and prioritization of important CI coefficients, and (iv) The division of the full QUBO into smaller, independent sub-QUBOs. These sub-QUBOs can be annealed in parallel and subsequently merged for post-processing to recover the final configuration.
The calculated PDM values for molecules like MgF, CaF, SrF, and BaF exhibit excellent agreement with classical benchmark methods, including FCI and RCCSD results, with deviations often below 1%. The results are in agreement with those obtained from SA. The accuracy of the QAE algorithm is rooted in its ability to determine ground-state energies with remarkable precision, as shown in Tables 5 and 6, where absolute errors relative to FCI are consistently on the order of $10^{-5}\%$ or less. The analysis also shows a few anomalies in the PDM calculation. For the BeF molecule in the larger (14o, 7e) active space, QAE performs less accurately with some basis sets, possibly due to how the molecule’s electronic structure interacts with the sub-QUBO partitioning. Aside from these few cases, the method works well across different molecules and basis sets.
Importantly, this work represents the first demonstration of molecular property calculations using quantum annealers. Our findings indicate that QAE, when combined with finite-field methods and perturbative many electron basis selection, can serve as a reliable tool for evaluating electronic properties of molecules. Looking ahead, scaling up QAE to larger active spaces presents challenges due to increased encoding complexity and convergence [37, 38]. However, recent advances in quantum-classical workflows can treat an individual orbital as a qubit for example, varialtional quantum annealing(varQA) [39]. Future advances in quantum annealing hardware, such as new generation D-Wave Advantage2 QPU with Zephyr topology [35] and non-linear quadratic solver [40], may further reduce embedding overheads and improve solution accuracy. These developments are expected to enhance both the QUBO encoding and annealing stages of QAE. We believe that, with continued improvements in hardware and algorithm design—including better preprocessing and error mitigation techniques [34, 41]—quantum annealers may soon play a significant role in atomic and molecular computations, enabling studies beyond the capabilities of classical methods.
@@ Acknowledgements.
Appendix A Appendix A: Ground-State Energy Benchmark Data
@@
[Uncaptioned image] [Uncaptioned image]
@@
[Uncaptioned image] [Uncaptioned image]

### v2

@@ Abstract
In summary, we have demonstrated the efficacy of QAE algorithm in computing the PDMs of molecules. Unlike the traditional applications of QAE, which focus on finding the ground-state energy, our approach integrates perturbative techniques and the FFM to enable the calculation of molecular properties. This novel framework holds promise for a broad range of many-body physics applications. Key contributions of our work include the integration of the finite-field approach into the QAE workflow and the division of the full QUBO into smaller, independent sub-QUBOs by selecting important expansion coefficients. These sub-QUBOs can be annealed in parallel and subsequently merged for post-processing to recover the final configuration.
The calculated PDM values for MgF, CaF, SrF, and BaF exhibit excellent agreement with classical benchmark methods, including FCI and RCCSD results, with deviations of average below 1%. The results are in agreement with those obtained from SA. An important feature of relativistic coupled-cluster method is that it includes physical effects to all orders of perturbation in the residual Coulomb interaction for all n particle-n hole excitations. Tables I and II of supplementary material indicate that absolute errors relative to FCI are consistently on the order of $10^{-5}\%$ or less. The analysis also shows a few anomalies in the PDM calculation. For the BeF molecule in the larger (14o, 7e) active space, QAE performs less accurately with some basis sets, possibly due to how the molecule’s electronic structure interacts with the sub-QUBO partitioning. Aside from these few cases, the method works well across different molecules and basis sets.
Importantly, this work represents the first computation of a molecular property other than energy using a quantum annealer. Our findings indicate that QAE, when combined with finite-field methods and perturbative many-electron basis selection, could serve as a reliable tool for evaluating electronic properties of molecules. Scaling up QAE to larger active spaces presents challenges due to increased encoding complexity and subtle convergence issues [40, 41]. However, recent advances in quantum-classical workflows can treat an individual orbital as a qubit as in, variational quantum annealing (varQA) algorithm [42]. Future advances in quantum annealing hardware, such as new D-Wave Advantage2 QPU with Zephyr topology [36] and non-linear quadratic solver [43], may further reduce embedding overheads and improve solution accuracy. These developments are expected to enhance both the QUBO encoding and annealing stages of QAE. Quantum annealers may soon play a significant role in atomic and molecular computations, enabling studies beyond the capabilities of classical methods with continued improvements in hardware and algorithm design—including better preprocessing and error mitigation techniques [35, 44].
@@ Acknowledgements.
Appendix A SUPPLEMENTARY MATERIAL : Ground-State Energy Benchmark Data
@@
Refer to caption Refer to caption
@@
Refer to caption Refer to caption